# Conditional Distribution Compression via the Kernel Conditional Mean Embedding

### v3

@@ B.4.2 Statement and Proof of Lemma
$\left\langle\hat{\mu}_{Y\mid X=\bm{x}},\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}}\right\rangle_{\mathcal{H}_{l}}$ $=\left\langle\sum_{i,j=1}^{n}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\cdot),\sum_{s,t=1}^{m}k(\bm{x},\bm{x}^{\prime}_{s})W_{st}l(\bm{y}^{\prime}_{t},\cdot)\right\rangle_{\mathcal{H}_{l}}$
$\underset{\text{(a)}}{=}\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x},\bm{x}_{i})W_{ij}\left\langle l(\bm{y}_{j},\cdot),l(\bm{y}^{\prime}_{t},\cdot)\right\rangle_{\mathcal{H}_{l}}k(\bm{x},\bm{x}^{\prime}_{s})W_{st}$
$\underset{\text{(b)}}{=}\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\bm{y}^{\prime}_{t})k(\bm{x},\bm{x}^{\prime}_{s})W_{st}$
$\underset{\text{(c)}}{=}\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\bm{y}^{\prime}_{t})W_{ts}k(\bm{x}^{\prime}_{s},\bm{x})$
@@
$\mathbb{E}_{\bm{x}\sim\mathbb{P}_{X^{*}}}\left[\left\langle\hat{\mu}_{Y\mid X=\bm{x}},\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}}\right\rangle_{\mathcal{H}_{l}}\right]$ $=\mathbb{E}_{\bm{x}\sim\mathbb{P}_{X^{*}}}\left[\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\bm{y}^{\prime}_{t})W_{ts}k(\bm{x}^{\prime}_{s},\bm{x})\right]$
$\underset{\text{(a)}}{\approx}\sum_{r=1}^{q}\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x}^{*}_{r},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\bm{y}^{\prime}_{t})W_{ts}k(\bm{x}^{\prime}_{s},\bm{x}^{*}_{r})$
$=\text{Tr}\left(K_{\bm{X}^{*}\bm{X}}W_{\bm{X}\bm{X}}L_{\bm{Y}\bm{Y}^{\prime}}W_{\bm{X}^{\prime}\bm{X}^{\prime}}K_{\bm{X}^{\prime}\bm{X}^{*}}\right)$
where (a) follows from approximating the marginal expectation with the samples $\{\bm{x}^{*}_{i}\}_{i=1}^{p}\sim\mathbb{P}_{X^{*}}$ and (b) follows from $\text{Tr}(AB)=\sum_{i,i,j=1}^{n}a_{ij}b_{ji}$. Here we have defined $K_{\bm{X}^{*}\bm{X}}]_{ij}:=k(\bm{x}^{*}_{i},\bm{x}_{j})$, $W_{\bm{X}\bm{X}}:=\left(K_{\bm{X}\bm{X}}+\lambda I\right)^{-1}$, $L_{\bm{Y}\bm{Y}^{\prime}}:=[l(\bm{y}_{i},\bm{y}^{\prime}_{j}]_{ij}$, $W_{\bm{X}^{\prime}\bm{X}^{\prime}}:=\left(K_{\bm{X}^{\prime}\bm{X}^{\prime}}+\lambda I\right)^{-1}$, and $K_{\bm{X}^{\prime}\bm{X}^{*}}]_{ij}:=k(\bm{x}^{\prime}_{i},\bm{x}^{*}_{j})$.
@@
$\widehat{\text{AMCMD}^{2}}\left[\mathbb{P}_{X^{*}},\mathbb{P}_{Y\mid X},\mathbb{P}_{Y^{\prime}\mid X^{\prime}}\right]$ $:=\mathbb{E}_{\bm{x}\sim\mathbb{P}_{X^{*}}}\left[\|\hat{\mu}_{Y\mid X=\bm{x}}-\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}}\|_{\mathcal{H}_{l}}^{2}\right]$
$=\text{Tr}\left(K_{\bm{X}^{*}\bm{X}}W_{\bm{X}\bm{X}}L_{\bm{Y}\bm{Y}}W_{\bm{X}\bm{X}}K_{\bm{X}\bm{X}^{*}}\right)$
$\;\;\;\;\;-2\text{Tr}\left(K_{\bm{X}^{*}\bm{X}}W_{\bm{X}\bm{X}}L_{\bm{Y}\bm{Y}^{\prime}}W_{\bm{X}^{\prime}\bm{X}^{\prime}}K_{\bm{X}^{\prime}\bm{X}^{*}}\right)$
$\;\;\;\;\;+\text{Tr}\left(K_{\bm{X}^{*}\bm{X}^{\prime}}W_{\bm{X}^{\prime}\bm{X}^{\prime}}L_{\bm{Y}^{\prime}\bm{Y}^{\prime}}W_{\bm{X}^{\prime}\bm{X}^{\prime}}K_{\bm{X}^{\prime}\bm{X}^{*}}\right).$
@@
To state Corollary 4.4 in full context, we first introduce some additional definitions and notation. Let $\mathcal{L}(\mathcal{H}_{l})$ denote the space of bounded linear operators from $\mathcal{H}_{l}$ to itself. An operator-valued kernel $\Gamma:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{H}_{l})$ induces a vector-valued reproducing kernel Hilbert space (vvRKHS), denoted $\mathcal{H}_{\Gamma}$, which consists of functions $f:\mathcal{X}\to\mathcal{H}_{l}$ (see [38] for further details). The vvRKHS $\mathcal{H}_{\Gamma}$ is $C_{0}$ if $\mathcal{H}_{\Gamma}\subseteq C_{0}(\mathcal{X},\mathcal{H}_{l})$, the space of continuous functions from $\mathcal{X}$ to $\mathcal{H}_{l}$ that vanish at infinity. Moreover, the kernel $\Gamma$ is called $C_{0}$-universal if it is $C_{0}$ and $\mathcal{H}_{\Gamma}$ is dense in $L^{2}(\mathcal{X},\mathbb{P}_{X};\mathcal{H}_{l})$ for any probability measure $\mathbb{P}_{X}$ on $\mathcal{X}$. Denoting the identity operator on $\mathcal{H}_{l}$ by $\mathcal{I}_{\mathcal{H}_{l}}$, it is shown in [38] that the kernel $\Gamma(x,x\prime)=k(x,x\prime)\mathcal{I}_{\mathcal{H}_{l}}$ is $C_{0}$-universal whenever $k$ is a universal scalar kernel, such as the Gaussian or Laplacian kernel. Hence, in the statement of Corollary B.7, the assumption that $k(\cdot,\cdot)$ is universal implies that $\Gamma$ is $C_{0}$-universal, with $\mathcal{H}_{\Gamma}$ as described above.
Note that the specific choice of kernel $\Gamma(x,x\prime)=k(x,x\prime)\mathcal{I}_{\mathcal{H}_{l}}$ leads to the form of the KCME estimator given in equation (1) [20], which we adopt in this work. More generally, the corollary could be stated under the assumption that $\mathcal{H}_{\Gamma}$ is induced by an arbitrary $C_{0}$-universal operator-valued kernel $\Gamma$, thereby removing the requirement that $k$ be universal—following the more general formulation in Theorem 4.5 of [20]. However, as stated, this would yield a potentially different form of the KCME estimator than the one used in (1).
@@ B.9 Derivation of ACKH Objective and Gradients
$\underset{(\bm{x},\bm{y})\in\mathcal{X}\times\mathcal{Y}}{\arg\min}\;\;\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\|\tilde{\mu}^{\mathcal{C}^{m}}_{Y\mid X=\bm{x}^{\prime}}\right\|^{2}_{\mathcal{H}_{l}}\right]-2\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\langle\mu_{Y\mid X=\bm{x}^{\prime}},\tilde{\mu}^{\mathcal{C}^{m}}_{Y\mid X=\bm{x}^{\prime}}\right\rangle_{\mathcal{H}_{l}}\right].$ (24)
@@
$=\frac{1}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right)-\frac{2}{n}\text{Tr}\left(\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right),$
@@
$\underset{\text{(c)}}{=}\frac{1}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right)-\frac{2}{n}\text{Tr}\left(\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right),$
@@
$A(\tilde{K}_{m}(\bm{x})+\lambda I)=\bar{K}_{m}(\bm{x})^{T}\;\;\text{for}\;\;A\in\mathbb{R}^{n\times m}$
@@
$\nabla_{\bm{y}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})=\frac{1}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\nabla_{\bm{y}}\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right)$ (25)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\nabla_{\bm{y}}\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right),$
$\nabla_{\bm{x}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})$ $=\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right)$ (26)
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\nabla_{\bm{x}}\tilde{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\nabla_{\bm{x}}\tilde{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\nabla_{\bm{x}}\bar{K}_{m}(\bm{x})^{T}\right).$
@@
$\nabla_{\tilde{\bm{x}}_{m}}\tilde{K}_{m}(\tilde{\bm{x}}_{m})\in\mathbb{R}^{m\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{x}}_{m}}\tilde{K}_{m}(\tilde{\bm{x}}_{m})\right]_{ij}:=\frac{\partial k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})}{\partial\tilde{\bm{x}}_{m}}\in\mathbb{R}^{d}$
$\nabla_{\tilde{\bm{x}}_{m}}\bar{K}_{m}(\tilde{\bm{x}}_{m})\in\mathbb{R}^{n\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{x}}_{m}}\bar{K}_{m}(\tilde{\bm{x}}_{m})\right]_{ij}:=\frac{\partial k(\bm{x}_{i},\tilde{\bm{x}}_{j})}{\partial\tilde{\bm{x}}_{m}}\in\mathbb{R}^{d}$
$\nabla_{\tilde{\bm{y}}_{m}}\tilde{K}_{m}(\tilde{\bm{y}}_{m})\in\mathbb{R}^{m\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{y}}_{m}}\tilde{L}_{m}(\tilde{\bm{y}}_{m})\right]_{ij}:=\frac{\partial l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})}{\partial\tilde{\bm{y}}_{m}}\in\mathbb{R}^{p}$
$\nabla_{\tilde{\bm{y}}_{m}}\bar{L}_{m}(\tilde{\bm{y}}_{m})\in\mathbb{R}^{n\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{y}}_{m}}\bar{L}_{m}(\tilde{\bm{y}}_{m})\right]_{ij}:=\frac{\partial l(\bm{y}_{i},\tilde{\bm{y}}_{j})}{\partial\tilde{\bm{y}}_{m}}\in\mathbb{R}^{p}$
@@
$\nabla_{\bm{y}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})=\frac{1}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\nabla_{\bm{y}}\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{T}\right)$

### v4

@@ B.4.2 Statement and Proof of Lemma
$\left\langle\hat{\mu}_{Y\mid X=\bm{x}},\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}}\right\rangle_{\mathcal{H}_{l}}$ $=\left\langle\sum_{i,j=1}^{n}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\cdot),\sum_{s,t=1}^{m}k(\bm{x},\bm{x}^{\prime}_{s})W^{\prime}_{st}l(\bm{y}^{\prime}_{t},\cdot)\right\rangle_{\mathcal{H}_{l}}$
$\underset{\text{(a)}}{=}\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x},\bm{x}_{i})W_{ij}\left\langle l(\bm{y}_{j},\cdot),l(\bm{y}^{\prime}_{t},\cdot)\right\rangle_{\mathcal{H}_{l}}k(\bm{x},\bm{x}^{\prime}_{s})W^{\prime}_{st}$
$\underset{\text{(b)}}{=}\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\bm{y}^{\prime}_{t})k(\bm{x},\bm{x}^{\prime}_{s})W^{\prime}_{st}$
$\underset{\text{(c)}}{=}\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\bm{y}^{\prime}_{t})W^{\prime}_{ts}k(\bm{x}^{\prime}_{s},\bm{x})$
@@
$\frac{1}{q}\sum_{r=1}^{q}\left\langle\hat{\mu}_{Y\mid X=\bm{x}^{*}_{r}},\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}^{*}_{r}}\right\rangle_{\mathcal{H}_{l}}$ $=\frac{1}{q}\sum_{r=1}^{q}\sum_{i,j=1}^{n}\sum_{s,t=1}^{m}k(\bm{x}^{*}_{r},\bm{x}_{i})W_{ij}l(\bm{y}_{j},\bm{y}^{\prime}_{t})W^{\prime}_{ts}k(\bm{x}^{\prime}_{s},\bm{x}^{*}_{r})$
$=\frac{1}{q}\text{Tr}\left(K_{\bm{X}^{*}\bm{X}}W_{\bm{X}\bm{X}}L_{\bm{Y}\bm{Y}^{\prime}}W_{\bm{X}^{\prime}\bm{X}^{\prime}}K_{\bm{X}^{\prime}\bm{X}^{*}}\right)$
where the second line follows from $\text{Tr}(AB)=\sum_{i,j=1}^{n}a_{ij}b_{ji}$. Here we have defined $K_{\bm{X}^{*}\bm{X}}]_{ij}:=k(\bm{x}^{*}_{i},\bm{x}_{j})$, $W_{\bm{X}\bm{X}}:=\left(K_{\bm{X}\bm{X}}+\lambda I\right)^{-1}$, $L_{\bm{Y}\bm{Y}^{\prime}}:=[l(\bm{y}_{i},\bm{y}^{\prime}_{j}]_{ij}$, $W_{\bm{X}^{\prime}\bm{X}^{\prime}}:=\left(K_{\bm{X}^{\prime}\bm{X}^{\prime}}+\lambda I\right)^{-1}$, and $K_{\bm{X}^{\prime}\bm{X}^{*}}]_{ij}:=k(\bm{x}^{\prime}_{i},\bm{x}^{*}_{j})$.
@@
$\widehat{\text{AMCMD}^{2}}\left[\mathbb{P}_{X^{*}},\mathbb{P}_{Y\mid X},\mathbb{P}_{Y^{\prime}\mid X^{\prime}}\right]$ $:=\frac{1}{q}\sum_{i=1}^{q}\left\|\hat{\mu}_{Y\mid X=\bm{x}^{*}_{i}}-\hat{\mu}_{Y^{\prime}\mid X^{\prime}=\bm{x}^{*}_{i}}\right\|_{\mathcal{H}_{l}}^{2}$
$=\frac{1}{q}\text{Tr}\left(K_{\bm{X}^{*}\bm{X}}W_{\bm{X}\bm{X}}L_{\bm{Y}\bm{Y}}W_{\bm{X}\bm{X}}K_{\bm{X}\bm{X}^{*}}\right)$
$\;\;\;\;\;-\frac{2}{q}\text{Tr}\left(K_{\bm{X}^{*}\bm{X}}W_{\bm{X}\bm{X}}L_{\bm{Y}\bm{Y}^{\prime}}W_{\bm{X}^{\prime}\bm{X}^{\prime}}K_{\bm{X}^{\prime}\bm{X}^{*}}\right)$
$\;\;\;\;\;+\frac{1}{q}\text{Tr}\left(K_{\bm{X}^{*}\bm{X}^{\prime}}W_{\bm{X}^{\prime}\bm{X}^{\prime}}L_{\bm{Y}^{\prime}\bm{Y}^{\prime}}W_{\bm{X}^{\prime}\bm{X}^{\prime}}K_{\bm{X}^{\prime}\bm{X}^{*}}\right).$
@@
To state Corollary 4.4 in full context, we first introduce some additional definitions and notation. Let $\mathcal{L}(\mathcal{H}_{l})$ denote the space of bounded linear operators from $\mathcal{H}_{l}$ to itself. An operator-valued kernel $\Gamma:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{H}_{l})$ induces a vector-valued reproducing kernel Hilbert space (vvRKHS), denoted $\mathcal{H}_{\Gamma}$, which consists of functions $f:\mathcal{X}\to\mathcal{H}_{l}$ (see [12] for further details). The vvRKHS $\mathcal{H}_{\Gamma}$ is $C_{0}$ if $\mathcal{H}_{\Gamma}\subseteq C_{0}(\mathcal{X},\mathcal{H}_{l})$, the space of continuous functions from $\mathcal{X}$ to $\mathcal{H}_{l}$ that vanish at infinity. Moreover, the kernel $\Gamma$ is called $C_{0}$-universal if it is $C_{0}$ and $\mathcal{H}_{\Gamma}$ is dense in $L^{2}(\mathcal{X},\mathbb{P}_{X};\mathcal{H}_{l})$ for any probability measure $\mathbb{P}_{X}$ on $\mathcal{X}$. Denoting the identity operator on $\mathcal{H}_{l}$ by $\mathcal{I}_{\mathcal{H}_{l}}$, it is shown in [12] that the kernel $\Gamma(x,x^{\prime})=k(x,x^{\prime})\mathcal{I}_{\mathcal{H}_{l}}$ is $C_{0}$-universal whenever $k$ is a universal scalar kernel, such as the Gaussian or Laplacian kernel. Hence, in the statement of Corollary B.7, the assumption that $k(\cdot,\cdot)$ is universal implies that $\Gamma$ is $C_{0}$-universal, with $\mathcal{H}_{\Gamma}$ as described above.
Note that the specific choice of kernel $\Gamma(x,x^{\prime})=k(x,x^{\prime})\mathcal{I}_{\mathcal{H}_{l}}$ leads to the form of the KCME estimator given in equation (1) [59], which we adopt in this work. More generally, the corollary could be stated under the assumption that $\mathcal{H}_{\Gamma}$ is induced by an arbitrary $C_{0}$-universal operator-valued kernel $\Gamma$, thereby removing the requirement that $k$ be universal—following the more general formulation in Theorem 4.5 of [59]. However, as stated, this would yield a potentially different form of the KCME estimator than the one used in (1).
@@ B.9 Derivation of ACKH Objective and Gradients
$\underset{\bm{x},\bm{y}}{\arg\min}\;\;\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\|\tilde{\mu}^{\mathcal{C}^{m}}_{Y\mid X=\bm{x}^{\prime}}\right\|^{2}_{\mathcal{H}_{l}}\right]-2\mathbb{E}_{\bm{x}^{\prime}\sim\mathbb{P}_{X}}\left[\left\langle\mu_{Y\mid X=\bm{x}^{\prime}},\tilde{\mu}^{\mathcal{C}^{m}}_{Y\mid X=\bm{x}^{\prime}}\right\rangle_{\mathcal{H}_{l}}\right].$ (24)
@@
$=\frac{1}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right)-\frac{2}{n}\text{Tr}\left(\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right),$
@@
$\underset{\text{(c)}}{=}\frac{1}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right)-\frac{2}{n}\text{Tr}\left(\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right),$
@@
$A(\tilde{K}_{m}(\bm{x})+\lambda I)=\bar{K}_{m}(\bm{x})^{\top}\;\;\text{for}\;\;A\in\mathbb{R}^{n\times m}$
@@
$\nabla_{\bm{y}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})=\frac{1}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\nabla_{\bm{y}}\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right)$ (25)
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\nabla_{\bm{y}}\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right),$
$\nabla_{\bm{x}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})$ $=\frac{2}{n}\text{Tr}\left(\nabla_{\bm{x}}\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right)$ (26)
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\nabla_{\bm{x}}\tilde{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;+\frac{2}{n}\text{Tr}\left(\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\nabla_{\bm{x}}\tilde{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right)$
$\;\;\;\;\;\;-\frac{2}{n}\text{Tr}\left(\bar{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\nabla_{\bm{x}}\bar{K}_{m}(\bm{x})^{\top}\right).$
@@
$\nabla_{\tilde{\bm{x}}_{m}}\tilde{K}_{m}(\tilde{\bm{x}}_{m})\in\mathbb{R}^{m\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{x}}_{m}}\tilde{K}_{m}(\tilde{\bm{x}}_{m})\right]_{ij}:=\nabla_{\tilde{\bm{x}}_{m}}k(\tilde{\bm{x}}_{i},\tilde{\bm{x}}_{j})in\mathbb{R}^{d}$
$\nabla_{\tilde{\bm{x}}_{m}}\bar{K}_{m}(\tilde{\bm{x}}_{m})\in\mathbb{R}^{n\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{x}}_{m}}\bar{K}_{m}(\tilde{\bm{x}}_{m})\right]_{ij}:=\nabla_{\tilde{\bm{x}}_{m}}k(\bm{x}_{i},\tilde{\bm{x}}_{j})\in\mathbb{R}^{d}$
$\nabla_{\tilde{\bm{y}}_{m}}\tilde{K}_{m}(\tilde{\bm{y}}_{m})\in\mathbb{R}^{m\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{y}}_{m}}\tilde{L}_{m}(\tilde{\bm{y}}_{m})\right]_{ij}:=\nabla_{\tilde{\bm{y}}_{m}}l(\tilde{\bm{y}}_{i},\tilde{\bm{y}}_{j})\in\mathbb{R}^{p}$
$\nabla_{\tilde{\bm{y}}_{m}}\bar{L}_{m}(\tilde{\bm{y}}_{m})\in\mathbb{R}^{n\times m\times d},\;\;\text{with}\;\;\left[\nabla_{\tilde{\bm{y}}_{m}}\bar{L}_{m}(\tilde{\bm{y}}_{m})\right]_{ij}:=\nabla_{\tilde{\bm{y}}_{m}}l(\bm{y}_{i},\tilde{\bm{y}}_{j})\in\mathbb{R}^{p}$
@@
$\nabla_{\bm{y}}\mathcal{G}^{\mathcal{D}}_{m}(\bm{x},\bm{y})=\frac{1}{n}\text{Tr}\left(\bar{K}_{m}(\bm{x})\tilde{W}_{m}(\bm{x})\nabla_{\bm{y}}\tilde{L}_{m}(\bm{y})\tilde{W}_{m}(\bm{x})\bar{K}_{m}(\bm{x})^{\top}\right)$